# HueManity: Probing Fine-Grained Visual Perception in MLLMs

### v2

@@ 5 Results and Analysis
5.1 Human Performance: A Near-Perfect Baseline
Human volunteers established a crucial performance baseline, demonstrating exceptionally high accuracy and efficiency on the HueManity benchmark (detailed in Table 3). On the numerical recognition task, volunteers achieved a perfect 100% average accuracy, indicating that identifying two-digit strings within the dot patterns was straightforward. The more complex alphanumeric task also yielded a very strong average accuracy of 95.6%.
The minor errors in the alphanumeric task were insightful: annotator feedback consistently attributed confusion not to an inability to perceive the characters, but to difficulty differentiating visually similar character forms (e.g., ‘s’, ‘c’, ‘w’ upper and lower case variants). This distinction is important, as it underscores that the primary perceptual challenge of extracting patterns from the dots was readily overcome by the human visual system. Furthermore, volunteers processed these images with remarkable efficiency, typically requiring less than one second per image to identify the characters.
Near-perfect and rapid human scores offer crucial insights into MLLM performance on this task for several key reasons:
  • Confirms Task Solvability and Stimulus Clarity: High human accuracy confirms the stimuli provide unambiguous visual information for character identification, showing the tasks are well-defined and solvable by proficient visual systems.
  • Establishes a Performance Ceiling: Human performance sets a clear accuracy ceiling, highlighting the target for machine perception systems on this (for humans) perceptually simple task.
  • Highlights the Nature of MLLM Deficiencies: Significant MLLM deviations from human scores indicate machine-specific limitations, not task impossibility or unclear stimuli. This contrast helps determine if MLLM struggles lie in initial perceptual grouping or later recognition.
  • Contextualizes Machine Errors: Human ease and minor confusions (e.g., similar glyphs) provide valuable context for analyzing MLLM error types and severity, pinpointing where machine perception significantly diverges from human visual understanding.
In essence, the human baseline shows HueManity tests foundational visual skills that humans execute efficiently, despite MLLM challenges. This context is vital for appreciating the scale and nature of MLLM performance gaps.
5.2 ResNet50 Baseline: Demonstrating Task Learnability
To further contextualize the performance of Multimodal Large Language Models (MLLMs), we established a strong baseline using a fine-tuned ResNet50 He et al. (2015), a well-established traditional computer vision architecture. As detailed in Table 1 and Table 2, the ResNet50 model achieved high accuracy, scoring 96.5% on the numerical recognition task and 94.5% on the alphanumeric task.
This robust performance from a standard convolutional architecture, fine-tuned on a relatively small set of 2,000 images from the HueManity dataset (as described in Section 4.2), is highly informative. It clearly demonstrates that the task of identifying characters within the HueManity dot patterns is fundamentally learnable by established computer vision techniques.
Furthermore, achieving near-human accuracy with this setup suggests that the perceptual cues embedded in the stimuli are sufficiently rich and consistent for a focused model to learn effective recognition. This implies that the task, while designed to challenge nuanced perception, is not inherently intractable for AI and can be considered relatively easy for appropriately adapted vision models. The ResNet50 baseline’s success, therefore, pinpoints the challenges as MLLM-specific. This suggests that MLLM difficulties likely arise not from an inherently unsolvable visual task, but from how these larger, more general models currently process or prioritize fine-grained perceptual information.
Table 1: Accuracy on the number recognition task for human evaluators, ResNet50, and various MLLMs on both text masks and patterned HueManity images.
Mask Patterned Image
Humans (average) 100% 100%
ResNet50 - 96.5%
@@ 5.2 ResNet50 Baseline: Demonstrating Task Learnability
GPT-4.1 mini 100% 19.0%
GPT-4.1 100% 33.6%
Claude 3.7 Sonnet 100% 0.4%
@@
LLaVA-v1.6-7B 87.7% 3.3%
LLaVA-v1.6-13B 87.2% 8.1%
LLaVA-v1.6-34B 96.6% 7.8%
Mistral-small3.1-24b 100% 0.1%
Qwen VL Max 100% 0.2%
Pixtral 100% 1%
Table 2: Accuracy on the alphanumeric recognition task, comparing human performance with ResNet50 and MLLMs across text masks and patterned HueManity images.
Mask Patterned Image
Humans (average) 100% 95.6%
ResNet50 - 94.5%
API-based models
GPT-4.1 mini 72.4% 0.6%
GPT-4.1 80% 3.0%
Claude 3.7 Sonnet 82.2% 0%
Open-source models
LLaVA-v1.6-7B 15% 0%
LLaVA-v1.6-13B 31.8% 0.1%
LLaVA-v1.6-34B 27.1% 0%
Mistral-small3.1-24b 58.7% 0%
Qwen VL Max 83.5% 0%
Pixtral 65.8% 1.8%
Table 3: Human evaluation accuracies on 100-image subsets for both number and alphanumeric pattern recognition tasks.
Number Pattern Alphanumeric Pattern
Evaluator 1 100% 97%
Evaluator 2 100% 95%
Evaluator 3 100% 95%
Average 100% 95.6%
5.3 The Perceptual Gap: MLLM Performance vs. Baselines on HueManity
The performance of the nine evaluated Multimodal Large Language Models (MLLMs) on the HueManity benchmark stands in stark contrast to the near-perfect accuracies achieved by both human volunteers and the fine-tuned ResNet50 baseline He et al. (2015) (detailed in Table 1 and Table 2). Across both the numerical and alphanumeric tasks, all evaluated MLLMs exhibited a profound and consistent inability to match these baselines. For instance, even the best-performing MLLM achieved only 33.6% on the easier numeric task and a mere 3% on the harder alphanumeric task—figures that dramatically underscore their struggle.
This significant and consistent underperformance, on tasks demonstrably solvable by human vision and traditional AI methodologies, signals a critical gap in the fine-grained visual perception capabilities of current MLLMs. Such results necessitate a deeper exploration into the potential underlying reasons for these models’ difficulties.
Several characteristics inherent to the current design and training paradigms of many Multimodal Large Language Models (MLLMs) may contribute to their observed difficulties on tasks demanding nuanced visual perception. Firstly, the visual processing pipeline itself could be a factor. Vision encoders in MLLMs, though often powerful and pre-trained on diverse datasets (e.g., ViT variants Liu et al. (2023, 2024a); Agrawal et al. (2024); Bai et al. (2023a, 2025), are typically optimized for capturing semantic information and global scene context. This optimization focus might lead to the unintentional down-weighting or loss of extremely fine-grained local details—like subtle color shifts defining patterns within a dense, similarly featured background—that are crucial when strong semantic cues are absent. Furthermore, the interface used to connect these vision encoders to the language model—often a projection layer or a small MLP Liu et al. (2023, 2024a); Agrawal et al. (2024); Bai et al. (2023a, 2025)—necessarily transforms and condenses visual information. This critical step, while enabling multimodal fusion, might inadvertently act as an information bottleneck, potentially abstracting or losing the precise, high-resolution feature distinctions necessary for meticulous perceptual tasks.
Secondly, the nature of MLLM pre-training data and their resulting operational strengths may not fully align with the demands of certain fundamental perception challenges. MLLMs are typically pre-trained on vast web-scale corpora where images are often paired with textual descriptions or interleaved image-text documents Liu et al. (2023, 2024a); Agrawal et al. (2024); Bai et al. (2023a, 2025); Achiam et al. (2023). Such datasets, while fostering robust semantic alignment and impressive contextual understanding, might underrepresent stimuli requiring intensive perceptual organization purely from low-level visual features, without strong linguistic or clear object-based anchors. Consequently, MLLMs may not have adequately developed the specialized visual routines needed for tasks like consistently grouping spatially distributed elements based only on subtle shared properties (e.g., color similarity) within a field of similar distractors. Their celebrated success in many multimodal tasks often stems from leveraging their powerful integrated Large Language Models (LLMs) to interpret, reason about, and generate language from visual input Achiam et al. (2023); Anthropic (2025); Liu et al. (2023); Gong et al. (2023); Jiang et al. (2023). This reliance on higher-level, often text-mediated, understanding can be less effective when the core challenge demands direct, bottom-up visual processing and pattern extraction, rather than conceptual inference or semantic association. Such tasks require a foundational visual acuity that may not be a primary emergent outcome of training regimes largely focused on multimodal semantics and instruction following.
@@ 6 Conclusion and Future Directions
In this work, we introduced HueManity, a new benchmark dataset comprising 83,850 Ishihara-style images designed to rigorously evaluate the fine-grained visual perception of Multimodal Large Language Models (MLLMs). Our comprehensive evaluations of nine MLLMs revealed a stark performance gap: these advanced models struggled significantly (e.g.,  3-34% accuracy) on tasks readily solved by human volunteers (95-100% accuracy) and traditional computer vision models like a fine-tuned ResNet50 (94-96% accuracy). This highlights a critical limitation in current MLLM capabilities to discern patterns based on subtle visual cues in complex, noisy environments. We provide the dataset and generation code to the community to foster further research.
@@
The observed MLLM deficiencies likely stem from a combination of factors, including vision encoders not optimized for such fine-grained detail, information bottlenecks at the vision-language interface, pre-training data that underrepresents such perceptual challenges, and an over-reliance on text-mediated or high-level semantic reasoning for tasks demanding direct, bottom-up visual processing. Future work should therefore focus on several key directions to bridge this perceptual gap: 1) Developing novel MLLM architectures with improved visual front-ends and fusion mechanisms that better preserve and process low-level visual information. 2) Augmenting pre-training and fine-tuning datasets with stimuli specifically designed to enhance robust perception in visually complex scenarios. 3) Exploring training objectives that explicitly encourage the development of fundamental visual acuity and perceptual organization skills, independent of, yet complementary to, high-level reasoning. Addressing these areas will be crucial for advancing MLLMs towards more holistic and human-like visual understanding.
7 Limitations
While HueManity provides valuable insights into a specific facet of MLLM visual perception, we acknowledge certain limitations. Firstly, the benchmark evaluates a focused perceptual task: identifying two-character alphanumeric strings within Ishihara-style dot patterns. Although this serves as a robust proxy for the ability to discern patterns from visually complex backgrounds based on subtle cues, its direct generalization to the full spectrum of diverse, real-world fine-grained perceptual challenges MLLMs may encounter requires further investigation. Secondly, our current dataset, while substantial and systematically generated with controlled color contrasts, primarily explores variations in color and basic character forms. Future iterations or complementary benchmarks could expand to assess a wider array of perceptual dimensions, such as sensitivity to texture, orientation, motion, or more intricate compositions of embedded elements. Finally, while our study includes a diverse set of nine contemporary MLLMs, the MLLM landscape is evolving rapidly. Consequently, our findings represent a snapshot based on the models and their versions tested at the time of this study, and newer or differently architected models might exhibit different performance characteristics.
8 Usage of Generative AI tools
Appendix A A Brief Discussion on CIEDE2000 Color Difference

### v5

@@ 5 Results and Analysis
Table 2: Accuracy on the number and alphanumeric recognition tasks for human evaluators, ResNet50, and various MLLMs on both text masks and patterned HueManity images. $\pm$ denotes the Wilson Confidence intervals at 95% confidence.
Number Task Alphanumeric Task
Mask Pattern Mask Pattern
Random choice 1.11% 1.11% 0.029% 0.029%
Humans (average) - 99.37 $\pm$ 2.37% - 93.25 $\pm$ 5.08%
ResNet50 - 96.5 $\pm$ 1.15% - 94.5 $\pm$ 1.42%
@@
GPT-4.1 100 $\pm$ 0.19% 33.6 $\pm$ 2.92% 80 $\pm$ 2.47% 3.0 $\pm$ 1.07%
GPT-4.1 mini 100 $\pm$ 0.19% 19.0 $\pm$ 2.42% 72.4 $\pm$ 2.76% 0.6 $\pm$ 0.51%
Claude 3.7 Sonnet 100 $\pm$ 0.19% 0.4 $\pm$ 0.43% 82.2 $\pm$ 2.36% 0 $\pm$ 0.19%
@@
LLaVA-v1.6-34B 96.6 $\pm$ 1.13% 7.8 $\pm$ 1.66% 27.1 $\pm$ 2.75% 0 $\pm$ 0.19%
LLaVA-v1.6-13B 87.2 $\pm$ 2.07% 8.1 $\pm$ 1.69% 31.8 $\pm$ 2.88% 0.1 $\pm$ 0.27%
LLaVA-v1.6-7B 87.7 $\pm$ 2.03% 3.3 $\pm$ 1.11% 15 $\pm$ 2.21% 0 $\pm$ 0.19%
Mistral-small3.1-24b 100 $\pm$ 0.19% 0.1 $\pm$ 0.27% 58.7 $\pm$ 3.04% 0 $\pm$ 0.19%
Qwen VL Max 100 $\pm$ 0.19% 0.2 $\pm$ 0.33% 83.5 $\pm$ 2.29% 0 $\pm$ 0.19%
Pixtral 100 $\pm$ 0.19% 1 $\pm$ 0.64% 65.8 $\pm$ 2.93% 1.8 $\pm$ 0.84%
5.1 Human and ResNet50 Baseline Performance
To contextualize the performance of MLLMs, we established two critical baselines: human evaluators and a fine-tuned ResNet50 model. These baselines serve to confirm the task’s solvability and provide a clear performance ceiling.
Human evaluators demonstrated near-perfect accuracy, achieving 99.37% on the numerical task and 93.25% on the alphanumeric task. Annotator feedback indicated that the few errors on the alphanumeric task were due to confusion between visually similar characters (e.g., ‘s’ vs. ‘S’, ‘c’ vs. ‘C’, ‘w’ vs. ‘W’), not an inability to perceive the patterns. Annotators also report that the character recognition time was in the order of a few seconds. This near-flawless performance confirms that the stimuli are clear and the task is fundamentally solvable for a proficient visual system.
Similarly, a fine-tuned ResNet50 [21] model achieved high accuracy, scoring 96.5% on the numerical task and 94.5% on the alphanumeric task. This strong performance from a standard convolutional architecture, trained on only 2,000 examples, proves that the perceptual cues within the images are sufficient for established computer vision techniques to learn the task.
Together, these baselines establish that the HueManity benchmark is a well-posed and learnable perception challenge, suggesting that any failures by MLLMs are likely due to architectural or training-related limitations rather than the inherent difficulty of the task itself.
5.2 MLLM Performance
The performance of the nine evaluated MLLMs on HueManity sharply contrasts with the near-perfect accuracies of both humans and the ResNet50 baseline (Table 2). All MLLMs consistently struggled, with the best achieving only 33.6% on the numeric task and a mere 3% on the alphanumeric task.
As expected, most models demonstrated strong Optical Character Recognition (OCR) capabilities on the simple binary mask images. For the numeric mask task, nearly all models achieved perfect or near-perfect accuracy. Performance on the more complex alphanumeric mask was more varied, while API-based models like GPT-4.1 and Claude 3.7 Sonnet maintained high accuracy (>>80%), the open-source LLaVA family notably struggled (15-32% accuracy), indicating a weaker baseline OCR capability for complex character sets.
The models’ high OCR accuracy on masks makes their widespread failure on the patterned images even more striking. This drop in performance pinpoints the challenge as a perceptual failure, not a simple recognition one. On the “easy” numeric pattern task, a clear performance hierarchy emerged. GPT-4.1 was the only model with meaningful success at 33.6% accuracy. A second tier, including GPT-4.1 mini (19.0%) and the LLaVA family (3.3-8.1%), showed minimal capability. The remaining models, such as Claude 3.7 Sonnet and Qwen VL Max, failed almost completely, scoring near zero. On the “hard” alphanumeric pattern task, MLLM performance collapsed almost entirely. GPT-4.1 was again the top performer but with a mere 3.0% accuracy. It was followed by Pixtral at 1.8%. All other evaluated models scored less than 1%, demonstrating a near-total inability to perform perceptual grouping on this more complex task.
Several characteristics inherent to the current design and training paradigms of many MLLMs may contribute to their observed difficulties on tasks demanding nuanced visual perception.
Semantic Optimization in Pre-Training. MLLM vision encoders are typically optimized to capture high-level semantic information for tasks like scene understanding [29, 28, 2, 5, 6]. This focus on global context may cause the loss of fine-grained local details, such as the subtle color and texture cues that define the characters in HueManity. Furthermore, pre-training on web-scale datasets of images paired with descriptive text may not provide sufficient exposure to stimuli requiring intensive perceptual organization without strong semantic anchors.
Architectural Bottlenecks. The architectures themselves may contribute to this failure. Many MLLMs use Vision Transformers (ViTs) [16] that divide images into patches, a process that can disrupt fine-grained details within those patches. Additionally, the projection layers that connect the vision encoder to the language model can act as an information bottleneck, abstracting away precise, high-resolution feature distinctions that are essential for solving our task.
5.3 Can MLLMs Learn the Task? Probing with In-Context Learning and Fine-Tuning
In-Context Learning We tested whether providing in-context examples could improve performance. While this approach retained 100% accuracy for the simple “mask” images, it proved ineffective and often detrimental for the patterned Ishihara plates (Table 3). For most models, performance degraded as more examples were added. These results are consistent with [39]’s findings and we hypothesize that the visual complexity of the examples introduces noise that confuses the models, reinforcing the conclusion that the deficit is perceptual rather than a lack of contextual understanding.
Table 3: Few-shot performance on Ishihara Pattern for number task with 95% Wilson confidence intervals.
Model 0-shot (Baseline) 1-shot 3-shot 5-shot
GPT 4.1 33.6 $\pm$ 2.9% 34.0 $\pm$ 2.9% 35.0 $\pm$ 3.0% 31.0 $\pm$ 2.9%
GPT-4.1 mini 19.0 $\pm$ 2.5% 14.0 $\pm$ 2.2% 9.0 $\pm$ 1.8% 6.0 $\pm$ 1.5%
Pixtral Large 1.0 $\pm$ 0.6% 1.0 $\pm$ 0.6% 0.0 $\pm$ 0.2% 0.0 $\pm$ 0.2%
Claude 3.7 Sonnet 0.4 $\pm$ 0.4% 0.0 $\pm$ 0.2% 1.0 $\pm$ 0.6% 1.0 $\pm$ 0.6%
Qwen 2.5 VL Max 0.2 $\pm$ 0.3% 0.0 $\pm$ 0.2% 0.0 $\pm$ 0.2% 0.0 $\pm$ 0.2%
Table 4: Performance of fine-tuned Gemma3-4B and qualitative examples of predictions.
Training Data Task Accuracy Predictions for Ground Truth Values
ML sZ cT 1Q
N/A (Zero-Shot) Mask 80% ML SZ CT 1Q
Pattern 0% 76492385 739482561 74928536 749238561
Fine-Tuned on Masks Mask 94% ML sZ cT 1Q
Pattern 0% R1 M5 4model D4
Fine-Tuned on Ishihara Mask 0% 7s 24 n6 36
Pattern 0% tH vF 6t 58
MLLM Fine-tuning To determine if direct training could overcome the observed perceptual challenges, we LoRA fine-tuned the Gemma-3-4B model on 500 examples from the HueManity dataset using the HuggingFace SFTTrainer [^4] https://ai.google.dev/gemma/docs/core/huggingface_text_full_finetune with Adam optimizer, 2e-4 learning rate, and 3 epochs. The results of this limited experiment were striking and revealed a critical limitation. While fine-tuning on the simple text masks improved performance on that control task as expected, it provided no benefit for recognizing the patterned Ishihara images. More critically, fine-tuning on the patterned images not only failed to improve performance on the core task but also destroyed the model’s ability to solve the simple mask task, causing its accuracy to plummet from 80% to 0%. Instead of learning to perceive the characters, the model simply learned to output plausible-looking two-character strings (e.g., “4G”, “6t”) regardless of the visual input. This suggests the failure is rooted in deeper architectural limitations, as the model appears incapable of learning the necessary perceptual skill.
5.4 Ablations and Additional Analyses
Image Resolution Ablation. To understand the impact of image resolution, we conducted an ablation study by evaluating model performance across a range of resolutions from 300px to 1300px (Table 5). The results were twofold. First, models that performed poorly at the native 900px resolution, such as Qwen VL Max, failed across all tested resolutions, which reinforces the claim that they have a fundamental perceptual deficit. Second, we uncovered a “squinting effect” in the best-performing model, GPT-4.1, whose accuracy peaked at the lowest resolution, increasing from 34% to 49% at 300px. We hypothesize that downsampling acts as a low-pass filter, smoothing the high-frequency “noise” from the dot patterns and making the underlying character shapes more salient to the model.
Table 5: Image resolution ablation with 95% Wilson confidence intervals (N=100).
Model 300px 500px 700px 900px 1100px 1300px
GPT 4.1 $49.0\pm 9.8\%$ $29.0\pm 8.9\%$ $31.0\pm 9.1\%$ $34.0\pm 9.3\%$ $30.0\pm 9.0\%$ $29.2\pm 8.9\%$
GPT-4.1 mini $39.0\pm 9.6\%$ $18.2\pm 7.6\%$ $18.0\pm 7.6\%$ $18.0\pm 7.6\%$ $10.0\pm 5.9\%$ $6.0\pm 4.6\%$
Qwen VL Max $1.0\pm 2.0\%$ $1.0\pm 2.0\%$ $1.0\pm 2.0\%$ $1.0\pm 2.0\%$ $1.0\pm 2.0\%$ $1.0\pm 2.0\%$
Pixtral Large $1.0\pm 2.0\%$ $0.0\pm 1.9\%$ $1.0\pm 2.0\%$ $0\pm 1.9\% 0.0$$\pm 1.9\% 0.0$$0\pm 1.9\%$
MLLM Failure Patterns. A qualitative analysis of MLLM failures on the HueManity benchmark reveals that, unlike humans whose errors are predictable, MLLMs exhibit three distinct and more fundamental failure patterns when their perceptual abilities are overwhelmed (see Appendix E for more details).
  • $\circ$
    Hallucination: Models often generate completely unrelated and overly complex text, such as turning a two-character string into a full phrase (e.g., ‘MUST SEE’). This happens when the perceptual challenge overwhelms their visual processing.
  • $\circ$
    Evasion: They frequently engage in evasive behavior, either by describing the image as an ‘Ishihara test’ without attempting to identify the characters or by explicitly stating they are unable to perform the task. This suggests an internal confidence threshold triggers a pre-programmed failure response.
  • $\circ$
    Erratic Outputs: Their outputs are often erratic and unpredictable, ranging from random strings and nonsensical numbers to flawed, repetitive patterns (e.g., ‘[L1L1L1]’). This unpredictability points to a lack of robust and stable visual feature extraction.
Quantitative Correlation with Real-World Benchmarks. To validate that performance on HueManity translates to real-world capabilities, we correlated model rankings on our alphanumeric pattern task with the human-preference Elo ratings from Vision Arena [11], a benchmark for diverse, real-world multimodal tasks. Our analysis included the seven models common to both leaderboards: GPT-4.1, GPT-4.1 mini, Claude 3.7 Sonnet, Mistral Medium, Qwen VL Max, Pixtral Large, and LLaVA-v1.6-34B. The analysis revealed a strong and statistically significant positive correlation of $\bm{\rho=0.8214}$ ($\bm{p=0.0227}$). This result provides compelling empirical evidence that HueManity is not an isolated synthetic challenge but effectively measures a foundational perceptual skill that is highly predictive of an MLLM’s general performance on complex, real-world tasks. This validates our benchmark as a potent diagnostic tool for identifying critical and generalizable gaps in the core perceptual abilities of current models.
@@ 6 Conclusion and Future Directions
In this work, we introduce HueManity, a large-scale benchmark with 83,850 Ishihara-style images designed to probe the fine-grained visual perception of Multimodal Large Language Models (MLLMs). Our comprehensive evaluation of nine state-of-the-art models reveals a critical weakness: the best-performing MLLM achieved only (33.6%, 3%) accuracies on the numeric and alphanumeric Ishihara pattern recognition tasks, starkly contrasting with the near-perfect scores of human evaluators (99.37%, 93.25%) and a fine-tuned ResNet50 baseline (96.5%, 94.5%). We demonstrate that this perceptual deficit is not easily remedied, as neither in-context learning nor direct fine-tuning yielded improvements, suggesting the failure may be rooted in fundamental architectural limitations rather than the task’s inherent difficulty. The strong correlation ($\rho=0.82$) between HueManity performance and rankings on the real-world Vision Arena benchmark validates that our findings are not isolated to a synthetic task but are indicative of a broader, more generalizable weakness in MLLM perceptual abilities.
@@
The deficiencies observed likely stem from vision encoders optimized for high-level semantics at the expense of fine-grained detail, information bottlenecks at the vision-language interface, and pre-training datasets that lack sufficient perceptual challenges. To bridge this gap, future work should focus on developing novel MLLM architectures that preserve low-level visual information, augmenting training datasets with perceptually challenging stimuli, and designing training objectives that explicitly foster robust visual acuity independent of high-level semantic reasoning. We release our dataset and generation code to facilitate further research toward building more perceptually grounded MLLMs.
Appendix 0.A Model Inventory
Model Provider Version Temp / Top-p Max Tokens
GPT-4.1 mini OpenAI gpt-4.1-mini-2025-04-14 1.0 / 1.0 1024
GPT-4.1 OpenAI gpt-4.1-2025-04-14 1.0 / 1.0 1024
Claude 3.7 Sonnet Anthropic claude-3-7-sonnet-20250219 1.0 / 1.0 1024
LLaVA-v1.6-7B Ollama llava:7b 0.8 / 0.9 -1 (inf)
LLaVA-v1.6-13B Ollama llava:13b 0.8 / 0.9 -1 (inf)
LLaVA-v1.6-34B Ollama llava:34b 0.8 / 0.9 -1 (inf)
Mistral-small3.1-24b Mistral AI mistral-small-3.1-24b-instruct 0.7 / 1.0 1024
Qwen VL Max Qwen (OpenRouter) qwen/qwen-vl-max 1.0 / 1.0 1024
Pixtral Mistral AI pixtral-large-2411 0.7 / 1.0 1024
Table 6: Standardized evaluation settings per model. All models were evaluated between May 6-7, 2025. Ollama models were hosted on a single GPU machine with an NVIDIA A100. These generation values are the model/provider defaults and we stick to these values to avoid introducing noise to our evaluations.
Appendix 0.B Label Spaces & Chance Baselines
Character universe.
We define a character set $\mathcal{C}$ used to compose two-character labels. Starting from decimal digits and ASCII letters with case,
$\{0,\ldots,9,\,A,\ldots,Z,\,a,\ldots,z\}\quad(62\ \text{symbols}),$
we exclude four visually ambiguous characters: lowercase $\ell$ and uppercase $\{I,J,O\}$. Thus,
$|\mathcal{C}|=62-4=58.$
0.B.1 Tasks and exact label spaces
Numeric (two digits).
The numeric label space contains all two-digit numbers without a leading zero:
$\mathcal{N}=\{10,11,\ldots,99\},\qquad|\mathcal{N}|=90.$
A uniform random-guess baseline is therefore
$\Pr[\text{random correct}]=\frac{1}{|\mathcal{N}|}=\frac{1}{90}\approx 1.11\%.$
Alphanumeric (two characters).
The alphanumeric label space contains all ordered pairs from $\mathcal{C}$:
$\mathcal{A}=\mathcal{C}\times\mathcal{C},\qquad|\mathcal{A}|=58^{2}=3364.$
A uniform random-guess baseline is therefore
$\Pr[\text{random correct}]=\frac{1}{|\mathcal{A}|}=\frac{1}{3364}\approx 0.0297\%.$
0.B.2 Scoring policy
We compute case-sensitive exact-match accuracy over $\mathcal{N}$ or $\mathcal{A}$, as applicable. Normalization prior to scoring trims whitespace, removes a possible “Answer: " prefix, extracts the first two alphanumerics. Characters outside $\mathcal{C}$ are considered invalid.
Appendix 0.C A Brief Discussion on CIEDE2000 Color Difference